the table below pertains to iowan, an economy in which the typical consumer's basket consists of 4 pounds of pork and 3 bushles of corn.

year price of pork price of corn
2012 $20 per pound

$12 per bushel

2013 $25 per pound $18 per bushel

Refer to Table 24-3. If 2013 is the base year, then the CPI for 2013 was

a. 75.3

b. 100.0

c. 116.0

d. 132.8
Answer #1

Answer
option b

Base year basket cost =sum of (quantity of base year *price of base year)
Base year Basket cost =4*25+3*18
=154

CPI=(Current year basket cost /base year basket cost)*100
The current and base year basket cost is the same as the base year is 2013

CPI=(154/154)*100=100
the CPI is 100.0

Also, by default, a CPI in the base year is 100.